determine the shunt of a circuit if galvanometer has a scale of 10ma and a resistance of 3 ohm and current of 1a
1
Expert's answer
2020-08-12T16:18:16-0400
"i_GR_G=(i-i_G)R_s\\(0.01)(3)=(1-0.01)R_s\\R_s=0.03\Omega"